# Text-to-Scene with Large Reasoning Models

### v1

@@ Abstract
Prompt-driven scene synthesis allows users to generate complete 3D environments from textual descriptions. Current text-to-scene methods often struggle with complex geometries and object transformations, and tend to show weak adherence to complex instructions. We address these limitations by introducing Reason-3D, a text-to-scene model powered by large reasoning models (LRMs). Reason-3D integrates object retrieval using captions covering physical, functional, and contextual attributes. Reason-3D then places the selected objects based on implicit and explicit layout constraints, and refines their positions with collision-aware spatial reasoning. Evaluated on instructions ranging from simple to complex indoor configurations, Reason-3D significantly outperforms previous methods in human-rated visual fidelity, adherence to constraints, and asset retrieval quality. Beyond its contribution to the field of text-to-scene generation, our work showcases the advanced spatial reasoning abilities of modern LRMs. Additionally, we release the codebase to further the research in object retrieval and placement with LRMs. [^1] Demo & code: https://lucala.github.io/reason-3d-demo/
@@ 1 Introduction
The demand for customizable 3D scenes is rapidly growing across a wide range of fields, including interior design (Çelen et al. 2024; Fu et al. 2021; Dai et al. 2017), video game development (Raistrick et al. 2024; Yang et al. 2024c; Unity Technologies 2025), autonomous driving (Gao et al. ; Mao et al. 2024), robotics (Lee, Park, and Kim 2025; Bu et al. 2024; Po and Wetzstein 2024; Cen et al. 2024), embodied AI (Yang et al. 2024a; Lee, Park, and Kim 2025; Deitke et al. 2022; Yang et al. 2024c), and more (Wen et al. 2025). Generating plausible and physically‐feasible arrangements of objects from high- to low-level descriptions is therefore at the core of text-to-scene models. Traditionally, scene synthesis has been addressed by specialized models trained on annotated layout datasets or by engineered pipelines that encode spatial priors (Paschalidou et al. 2021; Tang et al. 2024). But these methods tend to constrain the ability of the model in some way, be it context confinement, visual fidelity, physical plausibility, or object transformation freedom. Recent advances in training strategies, such as Group Relative Policy Optimization (GRPO), have led to the emergence of Large Reasoning Models (LRMs)  (Shao et al. 2024). Unlike standard LLMs, LRMs are trained to perform multi-step reasoning by leveraging test-time compute through long reasoning traces, enabling them to reason about geometry, context, physical affordances, and object-function relationships (Xu et al. 2025). This shift enables context-free scene synthesis, where systems can generate and arrange 3D environments directly from natural language, without relying on predefined templates or training distributions. In this work, we present Reason-3D, a modular and highly flexible scene synthesis pipeline that leverages LRMs to construct indoor and outdoor 3D scenes from open-ended textual descriptions. Unlike prior approaches, Reason-3D requires no domain-specific training or architectural constraints. Given a scene prompt, our system extracts relevant objects using a combination of embedding-based retrieval and LLM-based semantic voting across three dimensions: physical, functional, and contextual relevance. We use a dual-stage placement process, where each object is placed in an optimal order autoregressively with the help of an LRM. Finally, Reason-3D refines outputs by making the LRM aware of potential collisions and resolving them. This architecture enables Reason-3D to operate entirely through natural language instructions. It supports complex spatial compositions, unusual object configurations, and even outdoor or hybrid environments without requiring any manual scripting or handcrafted scene rules. We evaluate Reason-3D on established indoor scene synthesis tasks where we significantly outperform prior methods in terms of plausibility and structural coherence (cf. Fig. 1). Furthermore, we highlight the generalization capabilities of Reason-3D by demonstrating applications in outdoor scenes and composition scenarios. Through an ablation study, we find that certain LRMs consistently outperform others across a diverse set of spatial reasoning tasks. We summarize our contributions as follows:
@@ 2 Related Work
Recent progress in indoor scene synthesis has primarily followed two distinct approaches. One line of work capitalizes on the strong generative capabilities of image-based models, often employing Neural Radiance Fields (NeRFs) or 3D Gaussian splats as output representations (Schult et al. 2024; Zhou et al. 2024; Epstein et al. 2024; Po and Wetzstein 2024). While these methods produce visually realistic results, the generated scenes lack object-level separability, making them unsuitable for downstream tasks requiring precise object manipulation or interaction. A second research direction focuses on structured scene generation using intermediate representations, such as scene graphs or layout templates, coupled with curated asset libraries to synthesize 3D environments with discrete objects (Feng et al. 2023; Yang et al. 2024c).
@@
Within this structured generation paradigm, learning-based generative models such as diffusion models have emerged as powerful tools for modeling spatial priors (Hu et al. 2024). DiffuScene (Tang et al. 2024) applies a denoising diffusion process to generate unordered sets of object attributes, synthesizing layouts from the 3D-FRONT dataset, which comprises 19,000 annotated indoor scenes (Fu et al. 2021). Other methods use model training with scene graphs (Yang et al. 2025; Zhai et al. 2024; Lin and Yadong 2023; Zhai et al. 2023) and without scene graphs (Paschalidou et al. 2021; Yang et al. 2024a; Ritchie, Wang, and Lin 2019). Despite their effectiveness, these models require task-specific training and often struggle to generalize beyond the distributions they were trained on. In contrast, our method performs scene synthesis in a zero-shot setting, relying purely on language-based reasoning without additional fine-tuning or supervision.
@@
The advent of Large Language Models (LLMs) has enabled open-vocabulary 3D scene synthesis, supporting the flexible generation of scenes without dependence on predefined labels or categories. These systems treat the LLM as a “design assistant” that outputs a high-level layout or scene graph, which is then materialized into 3D. Holodeck (Yang et al. 2024c) and others  (Çelen et al. 2024; Aguina-Kang et al. 2024) orchestrate multi-agent LLM systems to generate scene graphs or DSLs (Domain Specific Language), which are later optimized through layout engines. These prompt-based LLM planners are able to interpret open-ended descriptions but often require post-processing or optimization to convert text into numeric layouts, which is an inherent limitation. LayoutVLM (Sun et al. 2025) represents the latest advancement in object placement, employing VLMs to generate two mutually reinforcing representations from visually marked images, and a self-consistent decoding process to improve spatial planning.
@@
A straightforward approach to scene synthesis with LLMs involves directly querying the model for object positions and rotations. LayoutGPT (Feng et al. 2023) pioneered this strategy by retrieving relevant example layouts from a database to serve as in-context demonstrations. The GPT-based model then generates a new layout by predicting object bounding boxes conditioned on these exemplars. However, the reliance on a fixed library of example scenes limits generalization, as the diversity of layouts that can be produced is constrained by the database contents. Other approaches fine-tune pretrained LLMs for scene synthesis or editing (Bucher and Armeni 2025; Yang et al. 2024b), but similarly inherit dataset biases that hinder their generalizability. Concurrent work such as DirectLayout (Ran et al. 2025) prompts an LLM with chain-of-thought reasoning to generate a 2D bird’s-eye view layout, which is then lifted into 3D and refined. While promising, this method is fine-tuned on the 3D-FRONT dataset, again limiting its applicability outside the training distribution. Naively prompting standard LLMs that lack spatial multi-step reasoning capabilities to generate object coordinates often results in physically implausible outputs (e.g., overlapping objects, unrealistic placements, or violations of spatial constraints) because general-purpose non-reasoning LLMs do not inherently model geometry, scale, or collision dynamics. As a result, most recent LLM-based scene synthesis pipelines incorporate auxiliary mechanisms, such as example-based in-context retrieval, spatial reasoning decomposition, or fine-tuned scene grammar models, to compensate for limitations in raw layout prediction and ensure plausible, physically consistent scenes. These hybrid frameworks confirm that LLMs are well-suited for capturing high-level design semantics and following open-ended user instructions, but also underscore the limitations of purely text-based models when tasked with precise spatial generation, in the absence of reasoning steps. In contrast, our method operates entirely with off-the-shelf LLMs exhibiting multi-step spatial reasoning, also referred to as LRMs, requiring no task-specific fine-tuning. This enables generalization to more diverse and open-ended settings, including outdoor environments.
@@ 3 Methodology
Our proposed method, Reason-3D, consists of multiple stages (cf. Fig. 2). Broadly speaking, the pipeline is divided into two phases: object retrieval and object placement.
@@
The objects used in our pipeline are sourced from the Objaverse dataset (Deitke et al. 2023). To ensure reliable operation, these 3D objects must conform to a standardized structure: specifically, each object is required to be upright and oriented consistently along a canonical forward direction. While the system includes runtime mechanisms to detect and correct certain rotational misalignments, overall robustness is significantly improved when operating on a geometrically consistent dataset. Since objects imported from Objaverse do not always adhere to these conventions, we apply a preprocessing step involving a vision-language model (VLM) to analyze the object’s appearance and infer its intended orientation. The VLM receives four rendered views of the object and selects the image that most likely represents the front-facing orientation. The identified view is then used to reorient the object accordingly. The remaining discrepancies must be addressed dynamically during layout synthesis by the LRM, provided the object’s dimensions and context allow for such adjustments.
@@
Object placement and scene construction are carried out in two phases: initial placement and refinement, where Reason-3D leverages the multi-step reasoning capabilities of LRMs to sequentially determine each object’s position and rotation in the context of previously placed objects, with full control over all three axes of rotation. The model receives scene constraints, object metadata (name and size), as well as a list of current placements, but no visual feedback (only labeled bounding boxes), requiring it to reason purely from spatial metadata. Before placement, we address two key challenges. First, implicit spatial relations in the scene prompt are made explicit using an LRM extraction step. This involves querying the LRM to imagine the layout and giving as output object-specific constraints, both relative to other objects and absolute. Second, object dependencies are resolved via a placement priority list generated by another LRM, ensuring a logical ordering (e.g., placing a table before placing a plate on top of that table). During placement, the LRM relies on bounding box dimensions to determine spatial relationships. However, in dense scenes, collisions may occur because the model selectively considers only a subset of previously placed objects (typically those it deems locally relevant to the current placement). This limitation stems from constraints on reasoning depth and input length, which prevent the model from exhaustively evaluating all potential spatial conflicts across the entire scene. To mitigate these object collisions, we introduce a refinement step. After initial placement, the LRM receives detected collisions as input and revises object positions and rotations accordingly, processing each object in the same order as during initial placement. These collisions are overlaps of the objects’ bounding boxes. The detected collisions comprise a complete list of all overlapping bounding boxes with a small buffer. To alleviate the complexity of the geometric reasoning task, each object is also annotated with a ”size after rotation” attribute, which reflects the dimensions of the axis-aligned bounding box after applying the object’s rotation. We find that this precomputed property substantially improves the model’s ability to reason about spatial constraints in transformed coordinate spaces. Importantly, not all collisions are considered undesirable. Since reasoning is performed over bounding boxes, benign overlaps may be semantically valid (e.g., a trash bin under a table, even if the table and trash bin bounding boxes intersect). Therefore, the refinement step does not indiscriminately eliminate all collisions. Instead, the LRM evaluates each collision involving the current object being refined and determines whether it is contextually appropriate, allowing for fine-grained spatial reasoning and scene-specific tolerance of overlaps. This two-stage approach helps overcome common failure modes (e.g., wrongly ordered placements, unintended occlusions) and reduces reasoning load per step by injecting task-relevant geometric abstractions into the LRM’s input.
@@ 4 Results
In all experiments, we use the Gemini 2.5 model (Comanici et al. 2025) for object placement, object refinement, and image captioning (VLM). While we tested the pipeline on other LRMs, the results showed that Gemini 2.5 consistently yielded the best output quality. We use Gemini embedding-004 as the embedding model.
@@
We compare our system, Reason-3D, against two recent scene synthesis frameworks: Holodeck (Yang et al. 2024c) and LayoutVLM (Sun et al. 2025). Holodeck is a multi-agent system that uses LLMs to generate domain-specific layout programs, which are later optimized using layout engines. LayoutVLM, on the other hand, uses a VLM to produce visually grounded spatial representations and a self-consistency decoding mechanism to enhance spatial layout planning. We utilize the publicly released Objaverse subset provided by the LayoutVLM source code. This dataset is converted using Objathor (Yang et al. 2024c) for compatibility with Holodeck, while Reason-3D uses its own preprocessing. Since LayoutVLM does not support object retrieval, we supply it with objects retrieved using Reason-3D for the full scene synthesis experiments. For experiments focusing on layout design, the same set of objects is predetermined and supplied to both Reason-3D and LayoutVLM, alongside the textual prompts. In summary, we compare Reason-3D against Holodeck on the retrieval task, against LayoutVLM on layout generation, and against both systems in the full scene synthesis evaluation.
@@
We evaluate full scene synthesis using only a text prompt and room dimensions (cf. Fig. 3) for Reason-3D, LayoutVLM (Sun et al. 2025), and Holodeck (Yang et al. 2024c). LayoutVLM receives additional inputs, including a predefined object list and room configuration via its configuration file. In Table 1, we present the results of the human preference study using 60 participants. We find that Reason-3D produces more semantically aligned and spatially valid scenes compared to both baselines.
Models Win-rates [%] Elo $\uparrow$
HD LVLM R-3D
@@
Table 3: Results of the human evaluation on object placement. Each column indicates the instruction complexity (cf. Fig. 4). All results are below 0.2 standard deviation and on a scale from 1 (bad) to 5 (excellent). Overall, we observe that participants prefer Reason-3D over previous work, and that as the instruction complexity increases, Reason-3D can place objects more reasonably than previous work.
@@
To assess spatial reasoning capabilities in isolation, we benchmark the following LRMs: Gemini 2.5 Pro (Comanici et al. 2025), GPT-o3 (OpenAI 2025b), Claude Sonnet 4 (Anthropic 2025), DeepSeek-R1 (DeepSeek 2025). Additionally, we use GPT-4.1 (OpenAI 2025a) as a baseline due to its lack of test-time compute (i.e., no explicit multi-step reasoning). The evaluation covers three dimensions of spatial reasoning. First, we test the ability to reason in a transformed coordinate system, requiring the model to apply correct geometric transformations. Second, we assess whether the model can infer object orientation based on the dimensions of its axis-aligned bounding box. Finally, we combine both challenges in a complex spatial planning task, requiring integrated multi-object reasoning. We test the LRMs on three scenes (more information can be found in Appendix D). We conducted a human evaluation with 40 participants, where each participant rated a scene synthesized by two models in a head-to-head comparison. The results are presented in Table 4, with comparisons illustrated in Fig. 5. Furthermore, Table 4 summarizes the average token usage for each model, providing insight into the computational cost associated with different levels of reasoning.

### v2

@@ Abstract
Prompt-driven scene synthesis allows users to generate complete 3D environments from textual descriptions. Current text-to-scene methods often struggle with complex geometries and object transformations, and tend to show weak adherence to complex instructions. We address these limitations by introducing Reason-3D, a text-to-scene model powered by large reasoning models (LRMs). Reason-3D integrates object retrieval using captions covering physical, functional, and contextual attributes. Reason-3D then places the selected objects based on implicit and explicit layout constraints, and refines their positions with collision-aware spatial reasoning. Evaluated on instructions ranging from simple to complex indoor configurations, Reason-3D significantly outperforms previous methods in human-rated visual fidelity, adherence to constraints, and asset retrieval quality. Beyond its contribution to the field of text-to-scene generation, our work showcases the advanced spatial reasoning abilities of modern LRMs. Additionally, we release the codebase to further the research in object retrieval and placement with LRMs.
This is the extended version of the corresponding paper published in the AAAI 2026 proceedings.
Codehttps://github.com/ETH-DISCO/reason-3d
@@ 1 Introduction
The demand for customizable 3D scenes is rapidly growing across a wide range of fields, including interior design (ccelen2024design; fu20213d; dai2017scannet), video game development (raistrick2024infinigen; yang2024holodeck; unity_asset_store), autonomous driving (gaomagicdrive; mao2025dreamdrive), robotics (lee2025dynscene; bu2024closed; po2024compositional; cen2024using), embodied AI (yang2024physcene; lee2025dynscene; deitke2022procthor; yang2024holodeck), and more (wen20253dscenegenerationsurvey). Generating plausible and physically‐feasible arrangements of objects from high- to low-level descriptions is therefore at the core of text-to-scene models. Traditionally, scene synthesis has been addressed by specialized models trained on annotated layout datasets or by engineered pipelines that encode spatial priors (paschalidou2021atiss; tang2024diffuscene). But these methods tend to constrain the ability of the model in some way, be it context confinement, visual fidelity, physical plausibility, or object transformation freedom. Recent advances in training strategies, such as Group Relative Policy Optimization (GRPO), have led to the emergence of Large Reasoning Models (LRMs)  (shao2024deepseekmath). Unlike standard LLMs, LRMs are trained to perform multi-step reasoning by leveraging test-time compute through long reasoning traces, enabling them to reason about geometry, context, physical affordances, and object-function relationships (xu2025towards). This shift enables context-free scene synthesis, where systems can generate and arrange 3D environments directly from natural language, without relying on predefined templates or training distributions. In this work, we present Reason-3D, a modular and highly flexible scene synthesis pipeline that leverages LRMs to construct indoor and outdoor 3D scenes from open-ended textual descriptions. Unlike prior approaches, Reason-3D requires no domain-specific training or architectural constraints. Given a scene prompt, our system extracts relevant objects using a combination of embedding-based retrieval and LLM-based semantic voting across three dimensions: physical, functional, and contextual relevance. We use a dual-stage placement process, where each object is placed in an optimal order autoregressively with the help of an LRM. Finally, Reason-3D refines outputs by making the LRM aware of potential collisions and resolving them. This architecture enables Reason-3D to operate entirely through natural language instructions. It supports complex spatial compositions, unusual object configurations, and even outdoor or hybrid environments without requiring any manual scripting or handcrafted scene rules. We evaluate Reason-3D on established indoor scene synthesis tasks where we significantly outperform prior methods in terms of plausibility and structural coherence (see Fig. 1). Furthermore, we highlight the generalization capabilities of Reason-3D by demonstrating applications in outdoor scenes and composition scenarios. Through an ablation study, we find that certain LRMs consistently outperform others across a diverse set of spatial reasoning tasks. We summarize our contributions as follows:
@@ 2 Related Work
Recent progress in indoor scene synthesis has primarily followed two distinct approaches. One line of work capitalizes on the strong generative capabilities of image-based models, often employing Neural Radiance Fields (NeRFs) or 3D Gaussian splats as output representations (schult2024controlroom3d; zhou2024gala3d; epstein2024disentangled; po2024compositional). While these methods produce visually realistic results, the generated scenes lack object-level separability, making them unsuitable for downstream tasks requiring precise object manipulation or interaction. A second research direction focuses on structured scene generation using intermediate representations, such as scene graphs or layout templates, coupled with curated asset libraries to synthesize 3D environments with discrete objects (feng2023layoutgpt; yang2024holodeck).
@@
Within this structured generation paradigm, learning-based generative models such as diffusion models have emerged as powerful tools for modeling spatial priors (hu2024mixed). DiffuScene (tang2024diffuscene) applies a denoising diffusion process to generate unordered sets of object attributes, synthesizing layouts from the 3D-FRONT dataset, which comprises 19,000 annotated indoor scenes (fu20213d). Other methods use model training with scene graphs (yang2025mmgdreamer; zhai2024echoscene; lin2023instructscene; zhai2023commonscenes) and without scene graphs (paschalidou2021atiss; yang2024physcene; ritchie2019fast). Despite their effectiveness, these models require task-specific training and often struggle to generalize beyond the distributions they were trained on. In contrast, our method performs scene synthesis in a zero-shot setting, relying purely on language-based reasoning without additional fine-tuning or supervision.
@@
The advent of Large Language Models (LLMs) has enabled open-vocabulary 3D scene synthesis, supporting the flexible generation of scenes without dependence on predefined labels or categories. These systems treat the LLM as a “design assistant” that outputs a high-level layout or scene graph, which is then materialized into 3D. Holodeck (yang2024holodeck) and others  (ccelen2024design; aguina2024open) orchestrate multi-agent LLM systems to generate scene graphs or DSLs (Domain Specific Language), which are later optimized through layout engines. These prompt-based LLM planners are able to interpret open-ended descriptions but often require post-processing or optimization to convert text into numeric layouts, which is an inherent limitation. LayoutVLM (sun2025layoutvlm) represents the latest advancement in object placement, employing VLMs to generate two mutually reinforcing representations from visually marked images, and a self-consistent decoding process to improve spatial planning.
@@
A straightforward approach to scene synthesis with LLMs involves directly querying the model for object positions and rotations. LayoutGPT (feng2023layoutgpt) pioneered this strategy by retrieving relevant example layouts from a database to serve as in-context demonstrations. The GPT-based model then generates a new layout by predicting object bounding boxes conditioned on these exemplars. However, the reliance on a fixed library of example scenes limits generalization, as the diversity of layouts that can be produced is constrained by the database contents. Other approaches fine-tune pretrained LLMs for scene synthesis or editing (bucher2025respace; yang2024llplace), but similarly inherit dataset biases that hinder their generalizability. Concurrent work such as DirectLayout (ran2025direct) prompts an LLM with chain-of-thought reasoning to generate a 2D bird’s-eye view layout, which is then lifted into 3D and refined. While promising, this method is fine-tuned on the 3D-FRONT dataset, again limiting its applicability outside the training distribution. Naively prompting standard LLMs that lack spatial multi-step reasoning capabilities to generate object coordinates often results in physically implausible outputs (e.g., overlapping objects, unrealistic placements, or violations of spatial constraints) because general-purpose non-reasoning LLMs do not inherently model geometry, scale, or collision dynamics. As a result, most recent LLM-based scene synthesis pipelines incorporate auxiliary mechanisms, such as example-based in-context retrieval, spatial reasoning decomposition, or fine-tuned scene grammar models, to compensate for limitations in raw layout prediction and ensure plausible, physically consistent scenes. These hybrid frameworks confirm that LLMs are well-suited for capturing high-level design semantics and following open-ended user instructions, but also underscore the limitations of purely text-based models when tasked with precise spatial generation, in the absence of reasoning steps. In contrast, our method operates entirely with off-the-shelf LLMs exhibiting multi-step spatial reasoning, also referred to as LRMs, requiring no task-specific fine-tuning. This enables generalization to more diverse and open-ended settings, including outdoor environments.
@@ 3 Methodology
Our proposed method, Reason-3D, consists of multiple stages (see Fig. 2). Broadly speaking, the pipeline is divided into two phases: object retrieval and object placement.
@@
The objects used in our pipeline are sourced from the Objaverse dataset (objaverseXL). To ensure reliable operation, these 3D objects must conform to a standardized structure: specifically, each object is required to be upright and oriented consistently along a canonical forward direction. While the system includes runtime mechanisms to detect and correct certain rotational misalignments, overall robustness is significantly improved when operating on a geometrically consistent dataset. Since objects imported from Objaverse do not always adhere to these conventions, we apply a preprocessing step involving a vision-language model (VLM) to analyze the object’s appearance and infer its intended orientation. The VLM receives four rendered views of the object and selects the image that most likely represents the front-facing orientation (see Appendix A for the prompting strategy). The identified view is then used to reorient the object accordingly. The remaining discrepancies must be addressed dynamically during layout synthesis by the LRM, provided the object’s dimensions and context allow for such adjustments.
@@
Object placement and scene construction are carried out in two phases: initial placement and refinement, where Reason-3D leverages the multi-step reasoning of LRMs to sequentially determine each object’s position and rotation in the context of previously placed objects, with full control over all three axes of rotation. The model receives scene constraints, object metadata (name and size), as well as a list of current placements, but no visual feedback (only labeled bounding boxes), requiring it to reason purely from spatial metadata. Before placement, we address two key challenges. First, implicit spatial relations in the scene prompt are made explicit using an LRM extraction step. This involves querying the LRM to imagine the layout and giving as output object-specific constraints, both relative to other objects and absolute. Second, object dependencies are resolved via a placement priority list generated by another LRM, ensuring a logical ordering (e.g., placing a table before placing a plate on top of that table). During placement, the LRM relies on bounding box dimensions to determine spatial relationships. However, in dense scenes, collisions may occur because the model selectively considers only a subset of previously placed objects (typically those it deems locally relevant to the current placement). This limitation stems from constraints on reasoning depth and input length, which prevent the model from exhaustively evaluating all potential spatial conflicts across the entire scene. To mitigate these object collisions, we introduce a refinement step. After initial placement, the LRM receives detected collisions as input and revises object positions and rotations accordingly, processing each object in the same order as during initial placement. These collisions are overlaps of the objects’ bounding boxes. The detected collisions comprise a complete list of all overlapping bounding boxes with a small buffer. To alleviate the complexity of the geometric reasoning task, each object is also annotated with a “size after rotation” attribute, which reflects the dimensions of the axis-aligned bounding box after applying the object’s rotation. We find that this precomputed property substantially improves the model’s ability to reason about spatial constraints in transformed coordinate spaces. Importantly, not all collisions are considered undesirable. Since reasoning is performed over bounding boxes, benign overlaps may be semantically valid (e.g., a trash bin under a table, even if the table and trash bin bounding boxes intersect). Therefore, the refinement step does not indiscriminately eliminate all collisions. Instead, the LRM evaluates each collision involving the current object being refined and determines whether it is contextually appropriate, allowing for fine-grained spatial reasoning and scene-specific tolerance of overlaps. This two-stage approach helps overcome common failure modes (e.g., wrongly ordered placements, unintended occlusions) and reduces reasoning load by injecting task-relevant geometric abstractions into the LRM’s input.
@@ 4 Results
In all experiments, we use the Gemini 2.5 model (comanici2025gemini25pushingfrontier) for object placement, object refinement, and image captioning (VLM). While we tested the pipeline on other LRMs, the results showed that Gemini 2.5 consistently yielded the best output quality. We use Gemini embedding-004 as the embedding model.
@@
We compare our system, Reason-3D, against two recent scene synthesis frameworks: Holodeck (yang2024holodeck) and LayoutVLM (sun2025layoutvlm). Holodeck is a multi-agent system that uses LLMs to generate domain-specific layout programs, which are later optimized using layout engines. LayoutVLM, on the other hand, uses a VLM to produce visually grounded spatial representations and a self-consistency decoding mechanism to enhance spatial layout planning. We utilize the publicly released Objaverse subset provided by the LayoutVLM source code. This dataset is converted using Objathor (yang2024holodeck) for compatibility with Holodeck, while Reason-3D uses its own preprocessing. Since LayoutVLM does not support object retrieval, we supply it with objects retrieved using Reason-3D for the full scene synthesis experiments. For experiments focusing on layout design, the same set of objects is predetermined and supplied to both Reason-3D and LayoutVLM, alongside the textual prompts. In summary, we compare Reason-3D against Holodeck on the retrieval task, against LayoutVLM on layout generation, and against both systems in the full scene synthesis evaluation.
@@
We evaluate full scene synthesis using only a text prompt and room dimensions (see Fig. 3) for Reason-3D, LayoutVLM (sun2025layoutvlm), and Holodeck (yang2024holodeck). LayoutVLM receives additional inputs, including a predefined object list and room configuration via its configuration file. In Table 1, we present the results of the human preference study using 60 participants. We find that Reason-3D produces more semantically aligned and spatially valid scenes compared to both baselines.
Model Win-rates [%] Elo $\uparrow$
HD LVLM Reason-3D
@@
Table 3: Results of the human evaluation on object placement. Each column indicates the instruction complexity (see Fig. 4). All results are below 0.2 standard deviation and on a scale from 1 (bad) to 5 (excellent). Overall, we observe that participants prefer Reason-3D over previous work, and that as the instruction complexity increases, Reason-3D can place objects more reasonably than previous work.
@@
To assess spatial reasoning capabilities in isolation, we benchmark the following LRMs: Gemini 2.5 Pro (comanici2025gemini25pushingfrontier), GPT-o3 (openai2025o3_o4mini), Claude Sonnet 4 (anthropic2025claude4), DeepSeek-R1 (deepseek2025R1). Additionally, we use GPT-4.1 (openai2025gpt4_1) as a baseline due to its lack of test-time compute (i.e., no explicit multi-step reasoning). The evaluation covers three dimensions of spatial reasoning. First, we test the ability to reason in a transformed coordinate system, requiring the model to apply correct geometric transformations. Second, we assess whether the model can infer object orientation based on the dimensions of its axis-aligned bounding box. Finally, we combine both challenges in a complex spatial planning task, requiring integrated multi-object reasoning. We test the LRMs on three scenes (more information can be found in Appendix D). We conducted a human evaluation with 40 participants, where each participant rated a scene synthesized by two models in a head-to-head comparison. The results are presented in Table 4, with comparisons illustrated in Fig. 5. Furthermore, Table 4 summarizes the average token usage for each model, providing insight into the computational cost associated with different levels of reasoning.